# Grand-Canonical Typicality

### v1

@@ 6 Derivation of Statements 2a and 2b: Grand-Canonical Typicality
$u_{\mathbb{S}(\mathscr{H}_{R})}\Bigl\{\Psi\in\mathbb{S}(\mathscr{H}_{R}):\Bigl\|\operatorname{tr}^{B}|\Psi\rangle\langle\Psi|-\operatorname{tr}^{B}\hat{\rho}_{R}\Bigr\|_{\operatorname{tr}}\geq\eta+\frac{d^{S}}{\sqrt{d_{R}}}\Bigr\}\leq 4\exp\Big(-\frac{\eta^{2}d_{R}}{18\pi^{3}}\Bigr)$ (83)
@@
Applying this to $\mathscr{H}_{R}:=\mathscr{H}_{\mathrm{gmc}}$ with $B=S^{c}$ and using that $\dim\mathscr{H}_{\mathrm{gmc}}$ is large (see before (79)), we obtain that for most $\Psi\in\mathbb{S}(\mathscr{H}_{\mathrm{gmc}})$,
$\operatorname{tr}^{S^{c}}|\Psi\rangle\langle\Psi|\approx\operatorname{tr}^{S^{c}}\hat{\rho}_{\mathrm{gmc}}\,.$ (84)
By Statement 1a, $\operatorname{tr}^{S^{c}}\hat{\rho}_{\mathrm{gmc}}\approx\hat{\rho}^{S}_{\mathrm{gG}}$, which yields the desired result.
@@ 7.1 GAP Measure
Further investigation and discussion of GAP measures can be found in [16, Sec. 3], [48, 33, 14, 44, 50].
@@ 7.2 Derivation of Statement 3
We apply this to $B=S^{c}$. By Statement 2a, for most $\Psi\in\mathbb{S}(\mathscr{H}_{\mathrm{gmc}})$, $\operatorname{tr}^{B}|\Psi\rangle\langle\Psi|\approx\hat{\rho}_{\mathrm{gG}}^{S}$. By (91), [^5] Theorems 3 and 4 in [14] provide precise versions of this reasoning. $\mathrm{GAP}_{\operatorname{tr}^{B}|\Psi\rangle\langle\Psi|}\approx\mathrm{GAP}_{\hat{\rho}_{\mathrm{gG}}^{S}}$. Thus, $\mathscr{L}_{\psi^{S}}\approx\mathrm{GAP}_{\hat{\rho}_{\mathrm{gG}}^{S}}$ as claimed (and likewise in the setting of Statement 2b).
@@
Another reasoning leads to the related result that $\mathscr{L}_{\psi^{S}}$, when averaged over $\Psi\in\mathbb{S}(\mathscr{H}_{\mathrm{gmc}})$, is close to $\mathrm{GAP}_{\hat{\rho}_{\mathrm{gG}}^{S}}$. It is based on the property [16, Sec. 3, Property 3] that “GAP is hereditary,” i.e., that if $\Psi\in\mathbb{S}(\mathscr{H}^{S}\otimes\mathscr{H}^{B})$ has distribution $\mathrm{GAP}_{\hat{\rho}^{S}\otimes\hat{\rho}^{B}}$, then for any fixed ONB $b$ of $\mathscr{H}^{B}$, $\mathbb{E}_{\Psi}\mathscr{L}_{\psi^{S}}=\mathrm{GAP}_{\hat{\rho}^{S}}$. The difference to (94) is twofold: First, the equality is exact, not approximate; and second, it concerns the average over $\Psi$, while (94) is the case for most $\Psi$. For our application, the exactness does not help as it gets lost in other steps. The application here looks as follows: Suppose that $\hat{\rho}_{\mathrm{gmc}}\approx\hat{\rho}_{\mathrm{gG}}$. Then the typicality measure for $\Psi$ is $u_{\mathbb{S}(\mathscr{H}_{\mathrm{gmc}})}=\mathrm{GAP}_{\hat{\rho}_{\mathrm{gmc}}}\approx\mathrm{GAP}_{\hat{\rho}_{\mathrm{gG}}}\approx\mathrm{GAP}_{\hat{\rho}_{\mathrm{gG}}^{S}\otimes\hat{\rho}_{\mathrm{gG}}^{S^{c}}}$ by (39) and (91). By hereditarity, for any ONB $b$ of $\mathscr{H}^{S^{c}}$, $\mathbb{E}_{\Psi}\mathscr{L}_{\psi^{S}}\approx\mathrm{GAP}_{\hat{\rho}_{\mathrm{gG}}^{S}}$. $\diamond$
@@ 7.3 Derivation of Statement 4a
Since the derivation of Statement 4b is analogous, we focus on 4a.
An ONB diagonalizing certain commuting observables $\hat{A}_{1},\ldots,\hat{A}_{n}$ (such as $\hat{Q}_{1}^{S^{c}},\ldots,\hat{Q}_{K-1}^{S^{c}}$ or later $\hat{N}_{1}^{S^{c}},\ldots,\hat{N}_{r}^{S^{c}}$) consists of an ONB for each of their joint eigenspaces (which in the relevant cases will have high dimension). Thus, a purely random ONB diagonalizing $\hat{A}_{1},\ldots,\hat{A}_{n}$ consists of a purely random ONB in each of the joint eigenspaces. Since the construction of $\psi^{S}$ involves picking a random basis vector $b_{J}$, and since each basis vector lies in one of the joint eigenspaces, we ask what is the probability that the chosen basis vector lies in a particular joint eigenspace (say, the one with eigenvalues $\alpha_{1},\ldots,\alpha_{n}$, denoted $\mathscr{H}_{\alpha_{1}...\alpha_{n}}$). Since
@@
$\mathbb{P}\Bigl(b_{J}\in\mathscr{H}_{\alpha_{1}...\alpha_{n}}\Big|b\Bigr)$ $=\sum_{j:b_{j}\in\mathscr{H}_{\alpha_{1}...\alpha_{n}}}\mathbb{P}(J=j)$ (97a)
$=\Bigl\langle\Psi\Big|\Bigl(\hat{I}^{S}\otimes\sum_{j:b_{j}\in\mathscr{H}_{\alpha_{1}...\alpha_{n}}}|b_{j}\rangle\langle b_{j}|\Bigr)\Big|\Psi\Bigr\rangle$ (97b)
$=\langle\Psi|(\hat{I}^{S}\otimes\hat{P}_{\alpha_{1}...\alpha_{n}})|\Psi\rangle=:\tilde{p}(\alpha_{1},\ldots,\alpha_{n})\,$, (97c)
which is independent of the choice of basis $b$, as it depends only on the projection $\hat{P}_{\alpha_{1}...\alpha_{n}}$ to the joint eigenspace. As a consequence, the random variable $\psi^{S}$ can be constructed from a given $\Psi$ as follows: instead of first choosing a basis $b$ diagonalizing $\hat{A}_{1},\ldots,\hat{A}_{N}$ in all of $\mathscr{H}^{B}$ and then a random $J$ with Born distribution, we can just as well choose first a joint eigenspace $\mathscr{H}_{\alpha_{1}...\alpha_{n}}$ with Born distribution (97c), then a random ONB $b^{\prime}$ in this subspace, and then one basis vector $b^{\prime}_{J^{\prime}}$ with conditional Born distribution, given that the eigenspace was $\mathscr{H}_{\alpha_{1}...\alpha_{n}}$. Finally, we can form
@@
$\Psi^{\prime}=\frac{(\hat{I}^{S}\otimes\hat{P}_{\alpha_{1}...\alpha_{n}})\Psi}{\|(\hat{I}^{S}\otimes\hat{P}_{\alpha_{1}...\alpha_{n}})\Psi\|}\,$, (99)
which can be thought of as the collapsed wave function after a simultaneous quantum measurement of $\hat{A}_{1},\ldots,\hat{A}_{n}$ on $\Psi$ with outcomes $\alpha_{1},\ldots,\alpha_{n}$. Since $b^{\prime}_{J^{\prime}}\in\mathscr{H}_{\alpha_{1}...\alpha_{n}}$, $\psi^{S}$ can just as well be obtained from $\Psi^{\prime}$ instead of $\Psi$,
@@
with different normalizing constant $\mathcal{N}^{\prime}$. That is, after choosing $\alpha_{1},\ldots,\alpha_{n}$ with Born distribution, we can simply continue with $\Psi^{\prime}$, so that $\mathscr{H}^{B}$ gets replaced by $\mathscr{H}_{\alpha_{1}...\alpha_{n}}$. In particular, Theorem 2 of [14] (as described around (94) above) can be applied as long as $\dim\mathscr{H}_{\alpha_{1}...\alpha_{n}}\gg 1$, and yields that for most ONBs $b^{\prime}$ of $\mathscr{H}_{\alpha_{1}...\alpha_{n}}$, the distribution of $\psi^{S}$ is approximately $\mathrm{GAP}_{\operatorname{tr}^{B}|\Psi^{\prime}\rangle\langle\Psi^{\prime}|}$. Thus, including all possible values of $\alpha_{1},\ldots,\alpha_{n}$,
$\mathscr{L}_{\psi^{S}}\approx\bigoplus_{\alpha_{1}...\alpha_{n}}\tilde{p}(\alpha_{1},\ldots,\alpha_{n})~\mathrm{GAP}_{\operatorname{tr}^{B}|\Psi^{\prime}\rangle\langle\Psi^{\prime}|}\,.$ (101)
@@
$=\bigoplus_{q_{1}^{S}...q_{K-1}^{S}}\Biggl(\prod_{k=1}^{K-1}\mathds{1}_{\{q_{k}^{S}\}}(\hat{Q}_{k}^{S})\mathscr{H}^{S}\Biggr)\otimes\Biggl(\prod_{k=1}^{K-1}\mathds{1}_{\{Q_{k}-q_{k}^{S}\}}(\hat{Q}_{k}^{S^{c}})\mathscr{H}^{S^{c}}\Biggr)\,$, (102b)
@@
$\Bigl(\hat{P}(\hat{Q}_{1}^{S}=q_{1}^{S},\ldots)\otimes\hat{I}^{S^{c}}\Bigr)\Big|_{\widetilde{\mathscr{H}}}=\Bigl(\hat{I}^{S}\otimes\hat{P}(\hat{Q}_{1}^{S^{c}}=Q_{1}-q_{1}^{S},\ldots)\Bigr)\Big|_{\widetilde{\mathscr{H}}}\,$. (103)
@@
$\operatorname{tr}^{S^{c}}|\Psi^{\prime}\rangle\langle\Psi^{\prime}|$ $=(\mathcal{N}^{\prime})^{2}\operatorname{tr}^{S^{c}}\Biggl(\hat{P}(\hat{Q}_{1}^{S^{c}}=Q_{1}-q_{1}^{S},\ldots)~|\Psi\rangle\langle\Psi|~\hat{P}(\hat{Q}_{1}^{S^{c}}=Q_{1}-q_{1}^{S},\ldots)\Biggr)$ (105a)
$=(\mathcal{N}^{\prime})^{2}\operatorname{tr}^{S^{c}}\Biggl(\hat{P}(\hat{Q}_{1}^{S}=q_{1}^{S},\ldots)~|\Psi\rangle\langle\Psi|~\hat{P}(\hat{Q}_{1}^{S}=q_{1}^{S},\ldots)\Biggr)$ (105b)
$=(\mathcal{N}^{\prime})^{2}\hat{P}(\hat{Q}_{1}^{S}=q_{1}^{S},\ldots)~\operatorname{tr}^{S^{c}}\Bigl(|\Psi\rangle\langle\Psi|\Bigr)~\hat{P}(\hat{Q}_{1}^{S}=q_{1}^{S},\ldots)$ (105c)

### v2

@@ 6 Derivation of Statements 2a and 2b: Grand-Canonical Typicality
$u_{\mathbb{S}(\mathscr{H}_{R})}\left\{\Psi\in\mathbb{S}(\mathscr{H}_{R}):\Bigl\|\operatorname{tr}^{B}|\Psi\rangle\langle\Psi|-\operatorname{tr}^{B}\hat{\rho}_{R}\Bigr\|_{\operatorname{tr}}\geq\eta+\frac{d^{S}}{\sqrt{d_{R}}}\right\}\leq 4\exp\left(-\frac{\eta^{2}d_{R}}{18\pi^{3}}\right)$ (91)
@@
Applying this to $\mathscr{H}_{R}:=\mathscr{H}_{\mathrm{gmc}}$ with $B=\overline{S}$ and using that $\dim\mathscr{H}_{\mathrm{gmc}}$ is large, we obtain that for most $\Psi\in\mathbb{S}(\mathscr{H}_{\mathrm{gmc}})$,
$\operatorname{tr}^{\overline{S}}|\Psi\rangle\langle\Psi|\approx\operatorname{tr}^{\overline{S}}\hat{\rho}_{\mathrm{gmc}}\,.$ (92)
By Statement 1a, $\operatorname{tr}^{\overline{S}}\hat{\rho}_{\mathrm{gmc}}\approx\hat{\rho}^{S}_{\mathrm{gG}}$, which yields the desired result.
@@ 7.1 GAP Measure
Further investigation and discussion of GAP measures can be found in [19, Sec. 3], [54, 38, 17, 50, 56].
@@ 7.2 Derivation of Statement 3
We apply this to $B=\overline{S}$. By Statement 2a, for most $\Psi\in\mathbb{S}(\mathscr{H}_{\mathrm{gmc}})$, $\operatorname{tr}^{B}|\Psi\rangle\langle\Psi|\approx\hat{\rho}_{\mathrm{gG}}^{S}$. By (99), [^5] Theorems 3 and 4 in [17] provide precise versions of this reasoning. $\mathrm{GAP}_{\operatorname{tr}^{B}|\Psi\rangle\langle\Psi|}\approx\mathrm{GAP}_{\hat{\rho}_{\mathrm{gG}}^{S}}$. Thus, $\mathscr{L}_{\psi^{S}}\approx\mathrm{GAP}_{\hat{\rho}_{\mathrm{gG}}^{S}}$ as claimed (and likewise in the setting of Statement 2b).
@@
Another reasoning leads to the related result that $\mathscr{L}_{\psi^{S}}$, when averaged over $\Psi\in\mathbb{S}(\mathscr{H}_{\mathrm{gmc}})$, is close to $\mathrm{GAP}_{\hat{\rho}_{\mathrm{gG}}^{S}}$. It is based on the property [19, Sec. 3, Property 3] that “GAP is hereditary,” i.e., that if $\Psi\in\mathbb{S}(\mathscr{H}^{S}\otimes\mathscr{H}^{B})$ has distribution $\mathrm{GAP}_{\hat{\rho}^{S}\otimes\hat{\rho}^{B}}$, then for any fixed ONB $b$ of $\mathscr{H}^{B}$, $\mathbb{E}_{\Psi}\mathscr{L}_{\psi^{S}}=\mathrm{GAP}_{\hat{\rho}^{S}}$. The difference to (102) is twofold: First, the equality is exact, not approximate; and second, it concerns the average over $\Psi$, while (102) is the case for most $\Psi$. For our application, the exactness does not help as it gets lost in other steps. The application here looks as follows: Suppose that $\hat{\rho}_{\mathrm{gmc}}\approx\hat{\rho}_{\mathrm{gG}}$. Then the typicality measure for $\Psi$ is $u_{\mathbb{S}(\mathscr{H}_{\mathrm{gmc}})}=\mathrm{GAP}_{\hat{\rho}_{\mathrm{gmc}}}\approx\mathrm{GAP}_{\hat{\rho}_{\mathrm{gG}}}\approx\mathrm{GAP}_{\hat{\rho}_{\mathrm{gG}}^{S}\otimes\hat{\rho}_{\mathrm{gG}}^{\overline{S}}}$ by (40) and (99). By hereditarity, for any ONB $b$ of $\mathscr{H}^{\overline{S}}$, $\mathbb{E}_{\Psi}\mathscr{L}_{\psi^{S}}\approx\mathrm{GAP}_{\hat{\rho}_{\mathrm{gG}}^{S}}$. $\diamond$
@@ 7.3 Derivation of Statement 4a
An ONB diagonalizing certain commuting observables $\hat{A}_{1},\ldots,\hat{A}_{n}$ (such as $\hat{Q}_{1}^{\overline{S}},\ldots,\hat{Q}_{K-1}^{\overline{S}}$ or later $\hat{N}_{1}^{\overline{S}},\ldots,\hat{N}_{r}^{\overline{S}}$) consists of an ONB for each of their joint eigenspaces (which in the relevant cases will have high dimension). Thus, a purely random ONB diagonalizing $\hat{A}_{1},\ldots,\hat{A}_{n}$ consists of a purely random ONB in each of the joint eigenspaces. Since the construction of $\psi^{S}$ involves picking a random basis vector $b_{J}$, and since each basis vector lies in one of the joint eigenspaces, we ask what is the probability that the chosen basis vector lies in a particular joint eigenspace (say, the one with eigenvalues $a_{1},\ldots,a_{n}$, denoted $\mathscr{H}_{a_{1}...a_{n}}$). Since
@@
$\mathbb{P}\Bigl(b_{J}\in\mathscr{H}_{a_{1}...a_{n}}\Big|b\Bigr)$ $=\sum_{j:b_{j}\in\mathscr{H}_{a_{1}...a_{n}}}\mathbb{P}(J=j)$ (105a)
$=\Bigl\langle\Psi\Big|\Bigl(\hat{I}^{S}\otimes\!\!\!\sum_{j:b_{j}\in\mathscr{H}_{a_{1}...a_{n}}}\!\!\!|b_{j}\rangle\langle b_{j}|\Bigr)\Big|\Psi\Bigr\rangle$ (105b)
$=\langle\Psi|(\hat{I}^{S}\otimes\hat{P}_{a_{1}...a_{n}})|\Psi\rangle=:\tilde{p}(a_{1},\ldots,a_{n})\,$, (105c)
which is independent of the choice of basis $b$, as it depends only on the projection $\hat{P}_{a_{1}...a_{n}}$ to the joint eigenspace. As a consequence, the random variable $\psi^{S}$ can be constructed from a given $\Psi$ as follows: instead of first choosing a basis $b$ diagonalizing $\hat{A}_{1},\ldots,\hat{A}_{n}$ in all of $\mathscr{H}^{B}$ and then a random $J$ with Born distribution, we can just as well choose first a joint eigenspace $\mathscr{H}_{a_{1}...a_{n}}$ with Born distribution (105c), then a random ONB $b^{\prime}$ in this subspace, and then one basis vector $b^{\prime}_{J^{\prime}}$ with conditional Born distribution, given that the eigenspace was $\mathscr{H}_{a_{1}...a_{n}}$. Finally, we can form
@@
$\Psi^{\prime}=\frac{(\hat{I}^{S}\otimes\hat{P}_{a_{1}...a_{n}})\Psi}{\|(\hat{I}^{S}\otimes\hat{P}_{a_{1}...a_{n}})\Psi\|}\,$, (107)
which can be thought of as the collapsed wave function after a simultaneous quantum measurement of $\hat{A}_{1},\ldots,\hat{A}_{n}$ on $\Psi$ with outcomes $a_{1},\ldots,a_{n}$. Since $b^{\prime}_{J^{\prime}}\in\mathscr{H}_{a_{1}...a_{n}}$, $\psi^{S}$ can just as well be obtained from $\Psi^{\prime}$ instead of $\Psi$,
@@
with different normalizing constant $\mathcal{N}^{\prime}$. That is, after choosing $a_{1},\ldots,a_{n}$ with Born distribution, we can simply continue with $\Psi^{\prime}$, so that $\mathscr{H}^{B}$ gets replaced by $\mathscr{H}_{a_{1}...a_{n}}$. In particular, Theorem 2 of [17] (as described around (102) above) can be applied as long as $\dim\mathscr{H}_{a_{1}...a_{n}}\gg 1$, and yields that for most ONBs $b^{\prime}$ of $\mathscr{H}_{a_{1}...a_{n}}$, the distribution of $\psi^{S}$ is approximately $\mathrm{GAP}_{\operatorname{tr}^{B}|\Psi^{\prime}\rangle\langle\Psi^{\prime}|}$. Thus, including all possible values of $a_{1},\ldots,a_{n}$,
$\mathscr{L}_{\psi^{S}}\approx\bigoplus_{a_{1}...a_{n}}\tilde{p}(a_{1},\ldots,a_{n})~\mathrm{GAP}_{\operatorname{tr}^{B}|\Psi^{\prime}\rangle\langle\Psi^{\prime}|}\,.$ (109)
@@
$=\bigoplus_{q_{1}^{S}...q_{K-1}^{S}}\Biggl(\prod_{k=1}^{K-1}\mathds{1}_{\{q_{k}^{S}\}}(\hat{Q}_{k}^{S})\mathscr{H}^{S}\Biggr)\otimes\Biggl(\prod_{k=1}^{K-1}\mathds{1}_{\{Q_{k}-q_{k}^{S}\}}(\hat{Q}_{k}^{\overline{S}})\mathscr{H}^{\overline{S}}\Biggr)\,$, (110b)
@@
$\Bigl(\hat{P}(\hat{Q}_{1}^{S}=q_{1}^{S},\ldots)\otimes\hat{I}^{\overline{S}}\Bigr)\Big|_{\widetilde{\mathscr{H}}}=\Bigl(\hat{I}^{S}\otimes\hat{P}(\hat{Q}_{1}^{\overline{S}}=Q_{1}-q_{1}^{S},\ldots)\Bigr)\Big|_{\widetilde{\mathscr{H}}}\,$. (111)
@@
$\operatorname{tr}^{\overline{S}}|\Psi^{\prime}\rangle\langle\Psi^{\prime}|$ $=(\mathcal{N}^{\prime})^{2}\operatorname{tr}^{\overline{S}}\Biggl(\hat{P}(\hat{Q}_{1}^{\overline{S}}=Q_{1}-q_{1}^{S},\ldots)~|\Psi\rangle\langle\Psi|~\hat{P}(\hat{Q}_{1}^{\overline{S}}=Q_{1}-q_{1}^{S},\ldots)\Biggr)$ (113a)
$=(\mathcal{N}^{\prime})^{2}\operatorname{tr}^{\overline{S}}\Biggl(\hat{P}(\hat{Q}_{1}^{S}=q_{1}^{S},\ldots)~|\Psi\rangle\langle\Psi|~\hat{P}(\hat{Q}_{1}^{S}=q_{1}^{S},\ldots)\Biggr)$ (113b)
$=(\mathcal{N}^{\prime})^{2}\hat{P}(\hat{Q}_{1}^{S}=q_{1}^{S},\ldots)~\operatorname{tr}^{\overline{S}}\Bigl(|\Psi\rangle\langle\Psi|\Bigr)~\hat{P}(\hat{Q}_{1}^{S}=q_{1}^{S},\ldots)$ (113c)
@@
7.4 Derivation of Statement 4b
We apply (109) to $n=r$ and $\hat{A}_{i}=\hat{N}_{i}^{\overline{S}}$. By Theorem 1 in [49] for $t=0$, for most $\Psi\in\mathbb{S}(\mathscr{H}_{\mathrm{gmc}})$, $\tilde{p}(n_{1}^{\overline{S}},\ldots,n_{r}^{\overline{S}})\approx\operatorname{tr}(\hat{\rho}_{\mathrm{gmc}}\hat{P}_{n})=p_{n}(n_{1}^{\overline{S}},\ldots,n_{r}^{\overline{S}})$. Given $n_{1}^{\overline{S}},\ldots,n_{r}^{\overline{S}}$, $\Psi^{\prime}$ is GAP-distributed with density matrix $\hat{P}_{n}\hat{\rho}_{\mathrm{gmc}}\hat{P}_{n}/p_{n}(n_{1}^{\overline{S}},\ldots,n_{r}^{\overline{S}})$. (This can be shown.) By the results of [50], $\operatorname{tr}^{\overline{S}}|\Psi^{\prime}\rangle\langle\Psi^{\prime}|\approx\hat{\rho}(n_{1}^{\overline{S}},\ldots,n_{r}^{\overline{S}})$ with probability close to 1. Thus, (109) takes the form (59). This completes the derivation of Statement 4b.
Remark 5 can now be obtained as follows: When acting on $\mathscr{H}_{\mathrm{gmc}}$, $\hat{I}^{S}\otimes\hat{P}_{n}$ can be replaced by $\hat{P}\otimes\hat{P}_{n}$ with $\hat{P}$ as in (57) and $q_{k}^{S}:=Q_{k}-F_{k}(n_{1}^{\overline{S}},\ldots,n_{r}^{\overline{S}})$. Given that we can otherwise replace $\hat{\rho}_{\mathrm{gmc}}$ by $\hat{\rho}_{\mathrm{gG}}\approx\hat{\rho}_{\mathrm{gc}}$, and using (40), we find that
$p_{n}(n_{1}^{\overline{S}},\ldots,n_{r}^{\overline{S}})$ $=\operatorname{tr}\bigl[\hat{\rho}_{\mathrm{gmc}}(\hat{I}^{S}\otimes\hat{P}_{n})\bigr]$ (114a)
$=\operatorname{tr}\bigl[\hat{\rho}_{\mathrm{gmc}}(\hat{P}\otimes\hat{P}_{n})\bigr]$ (114b)
$\approx\operatorname{tr}\bigl[\hat{\rho}_{\mathrm{gG}}(\hat{P}\otimes\hat{P}_{n})\bigr]$ (114c)
$=\operatorname{tr}(\hat{\rho}_{\mathrm{gG}}^{S}\hat{P})\,\operatorname{tr}(\hat{\rho}_{\mathrm{gG}}^{\overline{S}}\hat{P}_{n})$ (114d)
and in the same way $\hat{\rho}(n_{1}^{\overline{S}},\ldots,n_{r}^{\overline{S}})\approx\hat{\rho}^{S}(q_{1}^{S},\ldots,q_{K-1}^{S})$ as in (58). Thus, all summands in (59) with equal values of $q_{1}^{S},\ldots,q_{K-1}^{S}$ are multiples of the same distribution $\mathrm{GAP}_{\hat{\rho}^{S}(q_{1}^{S}...q_{K-1}^{S})}$, and thus can be combined into a single summand with weight
$\sum_{\begin{subarray}{c}n_{1}^{\overline{S}}...n_{r}^{\overline{S}}:\\ Q_{k}-F_{k}(n_{1}^{\overline{S}},\ldots,n_{r}^{\overline{S}})=q_{k}^{S}\end{subarray}}p_{n}(n_{1}^{\overline{S}},\ldots,n_{r}^{\overline{S}})\approx\operatorname{tr}(\hat{\rho}_{\mathrm{gG}}^{S}\hat{P})=p(q_{1}^{S},\ldots,q_{K-1}^{S})\,$, (115)
which leads to the equations of Remark 5.
Example 2.
As a simple example we consider the case of one particle species $(r=1)$. In this situation there is no notion of chemical equilibrium but only the one of spatial equilibrium.
Let $\mathscr{H}=\mathscr{H}^{S}\otimes\mathscr{H}^{\overline{S}}$ and let $\hat{Q}_{1}=\hat{N}$ be the particle number operator of the total system. Obviously, $\hat{N}$ is (exactly) extensive and thus we can decompose it as
$\hat{N}=\hat{N}^{S}\otimes\hat{I}^{\overline{S}}+\hat{I}^{S}\otimes\hat{N}^{\overline{S}},$ (116)
where $\hat{N}^{S}$ and $\hat{N}^{\overline{S}}$ are the particle number operators of $S$ and $\overline{S}$, with eigenvalues $n^{S}$ and $n^{\overline{S}}$, acting on $\mathscr{H}^{S}$ and $\mathscr{H}^{\overline{S}}$ respectively.
Let $\Delta N=0$ and let $N\in\mathbb{N}$ be an eigenvalue of $\hat{N}$, so $\mathscr{H}_{\mathrm{gmc}}$ is characterized by the (exact) total particle number and a micro-canonical energy interval. By Statement 4b, for most $\Psi\in\mathbb{S}(\mathscr{H}_{\mathrm{gmc}})$ and most ONBs $b$ of $\mathscr{H}^{\overline{S}}$ that diagonalize $\hat{N}^{\overline{S}}$,
$\mathscr{L}_{\psi^{S}}\approx\bigoplus_{n^{\overline{S}}=0}^{N}p_{n}(n^{\overline{S}})\,\mathrm{GAP}_{\hat{\rho}(n^{\overline{S}})},$ (117)
where
$p_{n}(n^{\overline{S}})=\operatorname{tr}(\hat{\rho}_{\mathrm{gmc}}\hat{P}_{n}),\qquad\hat{\rho}(n^{\overline{S}})=\frac{1}{p_{n}(n^{\overline{S}})}\operatorname{tr}^{\overline{S}}\left[\hat{P}_{n}\hat{\rho}_{\mathrm{gmc}}\hat{P}_{n}\right],$ (118)
with $\hat{P}_{n}=\hat{P}(\hat{N}^{\overline{S}}=n^{\overline{S}})$. These quantities can be evaluated more explicitly: with the help of (111) and Statement 1b we find that
$p_{n}(n^{\overline{S}})$ $=\operatorname{tr}\left(\left(\hat{P}(\hat{N}^{S}=N-n^{\overline{S}})\otimes\hat{I}^{\overline{S}}\right)\hat{\rho}_{\mathrm{gmc}}\right)$ (119a)
$=\operatorname{tr}^{S}\left(\hat{P}(\hat{N}^{S}=N-n^{\overline{S}})\operatorname{tr}^{\overline{S}}(\hat{\rho}_{\mathrm{gmc}})\right)$ (119b)
[…]
$=\frac{1}{Z_{\mathrm{gc}}^{S}}e^{\beta\mu(N-n^{\overline{S}})}\operatorname{tr}^{S}\left(\hat{P}(\hat{N}^{S}=N-n^{\overline{S}})e^{-\beta\hat{H}_{0}^{S}}\right),$ (119d)
where $\hat{H}_{0}^{S}$ is the part of $\hat{H}_{0}$ acting only on $S$ and $\mu=\mu_{01}$. Likewise,
$\hat{\rho}(n^{\overline{S}})$ $=\frac{1}{p_{n}(n^{\overline{S}})}\operatorname{tr}^{\overline{S}}\left(\hat{P}(\hat{N}^{\overline{S}}=n^{\overline{S}})\hat{\rho}_{\mathrm{gmc}}\right)$ (120a)
$=\frac{1}{p_{n}(n^{\overline{S}})}\hat{P}(\hat{N}^{S}=N-n^{\overline{S}})\Bigl(\operatorname{tr}^{\overline{S}}\hat{\rho}_{\mathrm{gmc}}\Bigr)\hat{P}(\hat{N}^{S}=N-n^{\overline{S}})$ (120b)
$\approx\frac{1}{Z_{\mathrm{gc}}^{S}}\frac{1}{p_{n}(n^{\overline{S}})}\hat{P}(\hat{N}^{S}=N-n^{\overline{S}})\,e^{-\beta(\hat{H}_{0}^{S}-\mu\hat{N}^{S})}\,\hat{P}(\hat{N}^{S}=N-n^{\overline{S}})$ (120c)
$=\frac{1}{Z_{\mathrm{gc}}^{S}p_{n}(n^{\overline{S}})}e^{\beta\mu(N-n^{\overline{S}})}\hat{P}(\hat{N}^{S}=N-n^{\overline{S}})\,e^{-\beta\hat{H}_{0}^{S}}\,\hat{P}(\hat{N}^{S}=N-n^{\overline{S}}).$ (120d)
Therefore, $\hat{\rho}(n^{\overline{S}})$ can be interpreted as a canonical density matrix of the system $S$ with $N-n^{\overline{S}}$ particles and thus (117) is a mixture of GAP measures with canonical density matrices corresponding to different particle numbers in $S$.
As a by-product, we observe that (63) reduces in this example to (119d) with $q_{1}^{S}=n^{S}=N-n^{\overline{S}}$ because $\hat{Q}_{1}=\hat{N}$, and (64) reduces to (120d). Thus, the calculation above amounts to an independent confirmation of Remark 5 in this case that does not assume the validity of the general Gibbs principle. $\diamond$